1. The algebraic expression - 16t2 + 75t represents the height (in feet) of a golf ball hit at an angle of 30 degrees to the horizontal after t seconds. Evaluate the expression for t = 0, 1, 2, 3, and 4 seconds and then describe what happens to the golf ball as time passes.
Answer by Theo(13342)

the following graph shows what happens.
in the graph, y represents the height of the ball and x represents the number of seconds from launch.



the ball will rise until it reaches a peak at 2.34 seconds and then it will start falling back to earth again.

the graph coordinates are in (x,y) format.
x is the number of seconds since launch.
y is the height.

for example, after 4 seconds, the height of the ball is 44 feet.

what you are looking at is a quadratic equation, with all the properties of a quadratic equation.

the standard form of a quadratic equation is y = ax^2 + bx + c
a is the coefficient of the x^2 term.
b is the coefficient of the x term.
c is the constant term.
the maximum height is at x = -b/(2a) = -75/(2*-16) = 2.34375.
when x = that, y = -16*x^2 + 75x = 87.890625.
the graph coordinates are rounded to 3 decimal digits.